(a) Compute the force constant or spring constant for a spring suspension system that is displaced 2.40 cm when a load of 100.0 kg is place upon this system. Use SI units.

(b) Use the force or spring constant from part (a), compute the potential energy stored in the spring when the spring is displaced 2.40 cm.
Answer #1

a)

F=KX=mg

=>K=mg/X =100*9.8/0.024

K=40833.33 N/m

b)

U=(1/2)KX2=(1/2)*40833.33*0.0242

U=11.76 J
Answer #2

a) F = k*x

m*g = k*x

==> k = m*g/x

= 100*9.8/0.024

= 40833.3 N/m

b) U = 0.5*k*x^2

= 11.76 J